# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ 1 Introduction
Recently, it has been appreciated that the Narain ensemble of Afkhami-Jeddi:2020ezh ; Maloney:2020nni emerges as a limit from the duality between a finite ensemble of Narain “code CFTs” and an abelian 3d Chern–Simons with a compact gauge group summed over bulk topologies Aharony:2023zit . In this paper, we discuss this duality in detail. Code CFTs are defined by even self-dual error correcting codes of length $n$ over ${\mathbb{F}}_{p}\times{\mathbb{F}}_{p}$, which emerge as possible non-anomalous maximal subgroups of 1-form symmetry of the bulk $(U(1)\times U(1))_{p}^{n}$ Chern–Simons theory Barbar:2023ncl . We consider boundary CFTs on arbitrary Riemann surfaces of genus $g$, and show that the holographic duality that equates “boundary ensemble average = sum over bulk topologies” originates from a topic within representation theory known as Howe duality, or the theta correspondence Howe1973Preprint ; Howe1979 ; Howe1989 ; Howe1989a .
For prime $p$, the bulk sum over topologies reduces to the sum over handlebodies, and can be expressed as a Poincaré series, which is a sum over all modular images of the vacuum character
@@
We will rigorously show that (1) holds for all genus, and with non-zero fugacities of $U(1)^{n}\times U(1)^{n}$ symmetry. Our proof does not rely on explicitly computing both sides. Instead, we will show that the RHS and LHS are both invariant under the action of two commuting groups, and, using Howe duality for finite fields, that there is a unique such invariant element. This implies that they are proportional.
We shall also discuss complementary pictures, reinterpreting them in our language. The Hilbert space of $(U(1)\times U(1))^{n}_{p}$ Chern–Simons theory on a genus $g$ Riemann surface is isomorphic to $g$ tensor copies of $2n$ level-$p$ qudits. This allows us to give an interpretation of the duality in terms of quantum information theory. Namely, both the path integral of a code CFT at the boundary and the path integral of the bulk theory on any individual handlebody topology are quantum stabilizer states, such that the holographic duality becomes an identity between different families of stabilizer states. We also explain that the ensemble of code CFTs discussed in this paper is the orbifold groupoid Gaiotto:2020iye of a given Narain theory, providing a SymTFT interpretation of the boundary ensemble.
@@ 1.1 Ensemble averaging in holography
Coleman Coleman:1988cy ; Coleman:1988tj already pointed out in the 1980s that the possible loss of quantum coherence due to wormholes in quantum gravity could be solved by considering an “ensemble.” These ideas have taken a much more concrete form in the context of AdS/CFT. This was first observed for JT gravity – a simple 2d model of gravity plus a dilaton Jackiw:1984je ; Teitelboim:1983ux – which was shown to be dual to a random ensemble of matrices, rather than to a system with a single Hamiltonian Saad:2019lba . In the same spirit, refs. Afkhami-Jeddi:2020ezh ; Maloney:2020nni showed that the partition function computed by averaging over the ensemble of all Narain CFTs matches (the perturbative part of) 3d Chern–Simons theory with non-compact group, summed over all possible handlebody geometries sharing the same boundary. This sum is properly implemented by summing over all of unique images of the vacuum character under the modular group, an observation that was anticipated by Maloney:2007ud , where it was noted that the sum over all classical gravity configuration plus their one-loop corrections is a sum over images of the Virasoro vacuum block. In the case of the Narain ensemble, the correct starting seed is the vacuum character of $U(1)^{n}\times U(1)^{n}$ chrial symmetry. The result, at genus 1, is
@@
where $\eta$ is Dedekind’s eta function. The summand with $\gamma=1$ gives the contribution of thermal AdS. The group of transformations in the sum is over the quotient by the subgroup $\Gamma$ which acts trivially on the seed. The RHS of (2) can essentially be taken as the definition of a real-analytic Eisenstein series, leading to
@@
In what follows, we will refer to the sum over modular images of the vacuum character as a Poincaré series. The vacuum character and its modular images can be interpreted as a path integral on different 3d i.e. solid tori, or genus-one handlebodies. Poincaré series have been known to arise in 3d quantum gravity for some time Dijkgraaf:2000fq ; Krasnov:2000zq ; Maloney:2007ud ; Castro:2011zq ; Keller:2014xba ; Jian:2019ubz . Following the work on averaging over Narain moduli space Maloney:2020nni ; Afkhami-Jeddi:2020ezh a number of generalizations and extensions for various Narain ensembles have been discussed, including adding fugacities Datta:2021ftn , considering rational points or discrete subsets in the Narain moduli space Raeymaekers:2021ypf ; Meruliya:2021utr ; Raeymaekers:2023ras ; Aharony:2023zit ; Ashwinkumar:2021kav ; Ashwinkumar:2023jtz ; Ashwinkumar:2023ctt , orbifolds of the Narain moduli space Benjamin:2021wzr ; Kames-King:2023fpa ; Forste:2024zjt , WZW models Dong:2021wot ; Meruliya:2021lul , and via top-down constructions in string theory Heckman:2021vzx .
@@ 1.2 Code CFTs
An additive code ${\mathcal{C}}$ is a collection of elements (called codewords) in some abelian group $\mathscr{D}$. The simplest case is binary codes, for which $\mathscr{D}=\{0,1\}^{n}$. Via “Construction A” of Leech and Sloane LeechSloane1971 , a binary code can be used to define a lattice by embedding it in $({\mathbb{Z}}/\sqrt{2})^{n}\subset\mathbb{R}^{n}$. When the code is doubly-even and self-dual (type II), the corresponding lattice will be even and self-dual, and can therefore be used to define a chiral CFT Dolan:1989kf ; Dolan:1994st . The torus partition function of a chiral CFT based on double-even self-dual code is given by
@@
Recently, two of us noted that this average is precisely of the “Poincaré series” form – up to an overall coefficient, the average enumerator polynomial is a sum over “modular transformations” generated by (6) acting on $x^{n}$ Henriksson:2022dml . The six terms in (7) represent the six unique images of $x^{n}$ under all possible combinations of $S$ and $T$. Furthermore Henriksson:2022dml noted that the combinations $\theta_{3}(2\tau)/\eta(\tau)$ and $\theta_{2}(2\tau)/\eta(\tau)$ are precisely the characters of $SU(2)_{1}$ chiral algebra, with the seed $x^{n}$ corresponding to $SU(2)_{1}^{n}$ vacuum character. This led to a conjecture that the average over corresponding code CFTs is holographically dual to $SU(2)_{1}^{n}$ Chern–Simons theory. [^2] The example here is for codes related to chiral theories with $(c,\bar{c})=(n,0)$. A more general average for codes relevant for $c=\bar{c}$ CFTs, was discussed for CFTs in Henriksson:2022dml , see also Bannai2020 ; Nebe2023 in the mathematical literature .
@@ 1.3 Bulk TQFT
The $U(1)^{n}\times U(1)^{n}$ 3d Chern–Simons is a topological theory. The Hilbert space ${\mathcal{H}}^{g}$ of this theory placed on a genus $g$ Riemann surface $\varSigma$ is finite dimensional, with the basis vectors being certain RCFT conformal blocks Witten:1988hf . In the abelian case, there is a continuous variety of boundary conditions, called embeddings in Barbar:2023ncl , and the corresponding blocks are not holomorphic Belov:2005ze ; Aharony:2023zit . A path integral of this theory on $\varSigma$ times an interval $[0,1]$ would calculate a transition amplitude between two states in ${\mathcal{H}}^{g}$. When the theory admits gapped (topological) boundary conditions Kapustin:2010hk ; Kaidi:2021gbs , imposing such a b.c. at one of the boundaries would define a state at the other boundary, the corresponding wavefunction being the partition function of a certain 2d CFT on $\varSigma$.
@@
The bulk Chern–Simons theory coupled to topological gravity by summing over all handlebody topologies gives rise to the Poincaré series, i.e. the sum over modular images of the vacuum character. In case of the genus-one boundary, the vacuum character is the path integral of the CS theory placed on thermal AdS, a handlebody geometry with one particular cycle topologically trivial in the bulk. All other handlebodies are related to thermal AdS by modular transformations. The resulting sum is the Hartle–Hawking wavefunction of “CS gravity.” In fact this definition would call for a sum over all bulk topologies, which is the prescription in the general case Dymarsky:2024frx , but in certain cases such as those discussed in this paper, this sum simplifies to include only handlebodies, or in fact a finite sum over equivalence classes of handlebodies – see section 4.1 for a discussion.
@@ 1.4 Howe duality
The central statement of Howe duality concerns a symplectic group $\mathcal{S}=Sp(2N,F)$ over a field $F$, and “dual pairs,” which are a pair of subgroups of $\mathcal{S}$ which are mutually centralizing and act reductively on a space $V$. The study of dual pairs was introduced by Howe Howe1973Preprint ; Howe1979 ; Howe1989 ; Howe1989a . Howe duality says that a certain representation of $\mathcal{S}$ has a particular structure when restricted to the pair of commuting subgroups. In the present case we will have $F={\mathbb{F}}_{p}$ – the field with $p$ elements – and the dual pair is formed by a group $O(n,n,F)$ which acts transitively on the set of codes, and the symplectic group $Sp(2g,F)$ of modular transformations/MacWilliams identities.
@@
For every $g$, each code $\cal C$ has an associated vector $\sum_{c_{i}\in\cal C}\psi_{c_{1},\ldots,c_{g}}$ in this space, which fully determines the genus-$g$ partition function of the corresponding code CFT. Individual terms in the Poincaré series are also vectors in this space. To connect with the picture above, the space $\mathcal{X}$ is isomorphic to the Hilbert space ${\mathcal{H}}^{g}$ of $(U(1)\times U(1))^{n}_{p}$ CS theory on a genus $g$ Riemann surface, and to the Hilbert space of $2gn$ qudits. The groups $Sp(2g,{\mathbb{F}}_{p})$ and $O^{+}(2n,{\mathbb{F}}_{p})=O(n,n,{\mathbb{F}}_{p})$ both act linearly on $\mathcal{X}$. The group actions commute and $O^{+}(2n,{\mathbb{F}}_{p})\times Sp(2g,{\mathbb{F}}_{p})$ forms a reductive dual pair inside a larger symplectic group $Sp(4ng,{\mathbb{F}}_{p})$, which also acts linearly on $\mathcal{X}$. It is straightforward to show that both the ensemble average and the Poincaré series are invariant under both groups.
@@ 2.1.1 $\boldsymbol{(U(1)\times U(1))_{k}}$ Chern–Simons theory on a torus
For simplicity let us start with the level-$k$$AB$ theory:” a single pair of gauge fields $A$ and $B$ on a 3d manifold $\mathcal{M}$ whose boundary is a torus, $\varSigma=\partial\mathcal{M}$. Within our general story this corresponds to the case of $n=g=1$. The 3d action is
@@
The quantization of this theory is reviewed in Aharony:2023zit following Bos:1989wa ; Elitzur:1989nr , see also Belov:2005ze . The first important element is the wavefunction, defined by a path integral on the handlebody (solid torus) with possible Wilson loop insertions, and fixed boundary conditions for the fields $A$ and $B$ at $\partial M$. For level-$k$$AB$ theory” placed on a torus, the Hilbert space ${\mathcal{H}}$ is a linear combination of $k^{2}$ wavefunctions, known as non-holomorphic blocks in Gukov:2004id ; Belov:2005ze (and codeword blocks in Aharony:2023zit ), given by
@@
The pure phase factor is necessary to make $W[a,b]$ invariant under the shifts of $a,b$ mod $k$,, which is confirmed by its action on the wavefunctions (14) Aharony:2023zit ,
@@
Figure 1: The two Wilson loops that wrap a genus-1 surface. The $\gamma_{a}$ loop in (a) is contractible in the bulk so it acts as the identity, while $\gamma_{b}$ in (b) is non-contractible so it acts non-trivially on the vacuum.
@@ 2.1.2 General case: $(U(1)\times U(1))^{n}_{k}$ theory on a genus-$g$ Riemann surface
We now consider the case of a CS theory with $2n$ gauge fields placed on $\mathcal{M}$ with a genus $g$ boundary $\varSigma=\partial\mathcal{M}$. We choose the standard basis of cycles $\gamma_{\kappa}$, $\kappa=1,2,\dots 2g$ on $\varSigma$ with the symplectic intersection
@@
where we have used $\Lambda$ here to denote the generating matrix of the lattice $\Lambda$, and star denotes complex-conjugate. The $g=1$ wavefunctions of this theory, parameterized by elements of the discriminant group $c\in{\mathscr{D}}=\Lambda^{*}/\Lambda=({\mathbb{Z}}_{k}\times{\mathbb{Z}}_{k})^{n}$ are
@@
We will need the higher-genus generalization of the wavefunctions, forming a basis in the Hilbert space ${\mathcal{H}}^{g}$ of the CS theory on $\varSigma$ of genus $g$. These are defined by $g$-tuples of vectors $(c_{1},\ldots,c_{g})$, $c_{i}\in{\mathbb{Z}}_{k}^{2n}$. From the bulk point of view these are the path integrals on a 3d handlebody geometry $\mathcal{M}$ ending on $\partial M=\varSigma$, with the Wilson line insertions of charges $c_{i}$ wrapping $g$ non-shrinkable cycles. The genus $g$ handlebody is characterized by a modular parameter $\Omega$, see e.g. Maloney:2020nni . Written explicitly, while taking $\xi=\bar{\xi}=0$, the wavefunctions are
@@
A genus-$g$ wavefunction is a sum over $g$-tuples of vectors $\vec{v}_{I}$, with the index $I$ running from $1,\ldots,g$. In addition, each $\vec{v}_{I}$ is an $2n$-dimensional vector in the lattice $\Lambda^{*}$. In (31), the (implicit) vector indices of $\vec{v}_{I}$ and $\vec{v}_{J}$ are contracted with $\eta$ or with each other. A fixed orthogonal matrix $\mathcal{O}\in O(n,n,\mathbb{R})$ specifies the “embedding” of codes into the Narain moduli space. The sum in (31) is performed over all integer-valued vectors $m_{I}$. The denominator $\Phi(\Omega)$ stands for the one-loop contribution of fluctuations of gauge fields in the bulk, generalizing $|\eta(\tau)|^{2}$ for $g>1$. This factor would be the same for any Chern-Simons theory with the same signature of matrix $K_{ij}$ Belavin:1986cy ; DHoker:1986eaw ; Zograf1987 ; Zograf1987b ; Sarnak1987 ; Krasnov:2000zq ; McIntyre:2004xs ; Gaberdiel:2010jf .
@@
[…]
Here $I$ runs from $1$ to $g$ indexing cycles, and index $j,j^{\prime}$ run from $1$ to $2n$, indexing gauge fields $A^{j}=\{A_{1},\ldots,A_{n},B_{1},\ldots,B_{n}\}$. As in the case of $g=n=1$, the phase factor in (33) is introduced to make it well defined for $a,b\in{\mathbb{Z}}_{k}^{2n}$.
@@
Generalizing the $g=1$ case of section 2.1.1, the “vacuum” state $|0\rangle:=\Psi_{0,\ldots,0}$ is the CS path integral on the handlebody with shrinkable cycles $\gamma_{a}^{I}$. It can be defined by requiring that any Wilson loop wrapping shrinkable cycle acts trivially,

### v2

@@ 1 Introduction
Recently, it has been appreciated that the Narain ensemble of Afkhami-Jeddi:2020ezh ; Maloney:2020nni emerges as a limit from the duality between a finite ensemble of Narain “code CFTs” and an abelian 3d Chern–Simons with a compact gauge group summed over bulk topologies Aharony:2023zit . In this paper, we discuss this duality in detail. Code CFTs are defined by even self-dual error correcting codes of length $n$ over ${\mathbb{Z}}_{k}\times{\mathbb{Z}}_{k}$, which emerge as possible non-anomalous maximal subgroups of 1-form symmetry of the bulk $(U(1)\times U(1))_{k}^{n}$ Chern–Simons theory Barbar:2023ncl . We consider boundary CFTs on arbitrary Riemann surfaces of genus $g$, and show that the holographic duality that equates “boundary ensemble average = sum over bulk topologies” originates from a topic within representation theory known as Howe duality, or the theta correspondence Howe1973Preprint ; Howe1979 ; Howe1989 ; Howe1989a .
For prime $k=p$, where ${\mathbb{Z}}_{p}={\mathbb{F}}_{p}$ is a field, the bulk sum over topologies reduces to the sum over handlebodies, and can be expressed as a Poincaré series, which is a sum over all modular images of the vacuum character
@@
We will rigorously show that (1) holds for all genus, and with non-zero fugacities of the $U(1)^{n}\times U(1)^{n}$ symmetry. Our proof does not rely on explicitly computing both sides. Instead, we will show that the RHS and LHS are both invariant under the action of two commuting groups, and, using Howe duality for finite fields, that there is a unique such invariant element. This implies that they are proportional.
We shall also discuss complementary pictures, reinterpreting them in our language. The Hilbert space of $(U(1)\times U(1))^{n}_{p}$ Chern–Simons theory on a genus-$g$ Riemann surface is isomorphic to $g$ tensor copies of $2n$ level-$p$ qudits. This allows us to give an interpretation of the duality in terms of quantum information theory. Namely, both the path integral of a code CFT at the boundary and the path integral of the bulk theory on any individual handlebody topology are quantum stabilizer states – hence the holographic duality becomes an identity between different families of stabilizer states. We also explain that the ensemble of code CFTs discussed in this paper is the orbifold groupoid Gaiotto:2020iye of a given Narain theory, providing a SymTFT interpretation of the boundary ensemble.
@@ 1.1 Ensemble averaging in holography
Coleman Coleman:1988cy ; Coleman:1988tj already pointed out in the 1980s that the possible loss of quantum coherence due to wormholes in quantum gravity could be solved by considering an “ensemble.” These ideas have taken a much more concrete form in the context of AdS/CFT. This was first observed for JT gravity – a simple 2d model of gravity plus a dilaton Jackiw:1984je ; Teitelboim:1983ux – which was shown to be dual to a random ensemble of matrices, rather than to a system with a single Hamiltonian Saad:2019lba . In the same spirit, refs. Afkhami-Jeddi:2020ezh ; Maloney:2020nni showed that the partition function computed by averaging over the ensemble of all Narain CFTs matches (the perturbative part of) 3d Chern–Simons theory with non-compact group, summed over all possible handlebody geometries sharing the same boundary. This sum is properly implemented by summing over all of unique images of the vacuum character under the modular group, an observation that was anticipated by Maloney:2007ud , where it was noted that the sum over all classical gravity configurations plus their one-loop corrections is a sum over images of the Virasoro vacuum block. In the case of the Narain ensemble, the correct starting seed is the vacuum character of $U(1)^{n}\times U(1)^{n}$ chiral symmetry. The result, at genus 1, is
@@
where $\eta$ is Dedekind’s eta function. The summand with $\gamma=1$ gives the contribution of thermal AdS. The quotient in the sum is over the subgroup $\Gamma$ which acts trivially on the seed. The RHS of (2) can essentially be taken as the definition of a real-analytic Eisenstein series, leading to
@@
In what follows, we will refer to the sum over modular images of the vacuum character as a Poincaré series. The vacuum character and its modular images can be interpreted as a path integral on different 3d spaces, namely solid tori, or genus-one handlebodies. Poincaré series have been known to arise in 3d quantum gravity for some time Dijkgraaf:2000fq ; Krasnov:2000zq ; Maloney:2007ud ; Castro:2011zq ; Keller:2014xba ; Jian:2019ubz . Following the work on averaging over Narain moduli space Maloney:2020nni ; Afkhami-Jeddi:2020ezh a number of generalizations and extensions for various Narain ensembles have been discussed, including adding fugacities Datta:2021ftn , considering rational points or discrete subsets in the Narain moduli space Raeymaekers:2021ypf ; Meruliya:2021utr ; Raeymaekers:2023ras ; Aharony:2023zit ; Ashwinkumar:2021kav ; Ashwinkumar:2023jtz ; Ashwinkumar:2023ctt , orbifolds of the Narain moduli space Benjamin:2021wzr ; Kames-King:2023fpa ; Forste:2024zjt , WZW models Dong:2021wot ; Meruliya:2021lul , and via top-down constructions in string theory Heckman:2021vzx .
@@ 1.2 Code CFTs
An additive code $\mathcal{C}$} is a collection of elements (called codewords) in some abelian group $\mathscr{D}$. The simplest case is binary codes, for which $\mathscr{D}=\{0,1\}^{n}$. Via “Construction A” of Leech and Sloane LeechSloane1971 , a binary code can be used to define a lattice by embedding it in $({\mathbb{Z}}/\sqrt{2})^{n}\subset\mathbb{R}^{n}$. When the code is doubly-even and self-dual (type II), the corresponding lattice will be even and self-dual, and can therefore be used to define a chiral CFT Dolan:1989kf ; Dolan:1994st . The torus partition function of a chiral CFT based on doubly-even self-dual code is given by
@@
Recently, two of us noted that this average is precisely of the “Poincaré series” form – up to an overall coefficient, the average enumerator polynomial is a sum over “modular transformations” generated by (6) acting on $x^{n}$ Henriksson:2022dml . The six terms in (7) represent the six unique images of $x^{n}$ under all possible combinations of $S$ and $T$. Furthermore, Henriksson:2022dml noted that the combinations $\theta_{3}(2\tau)/\eta(\tau)$ and $\theta_{2}(2\tau)/\eta(\tau)$ are precisely the characters of $SU(2)_{1}$ chiral algebra, with the seed $x^{n}$ corresponding to $SU(2)_{1}^{n}$ vacuum character. This led to a conjecture that the average over corresponding code CFTs is holographically dual to $SU(2)_{1}^{n}$ Chern–Simons theory. [^2] The example here is for codes related to chiral theories with $(c,\bar{c})=(n,0)$. A more general average for codes relevant for $c=\bar{c}$ CFTs, was discussed for CFTs in Henriksson:2022dml , see also Bannai2020 ; Nebe2023 in the mathematical literature
@@ 1.3 Bulk TQFT
The $U(1)^{n}\times U(1)^{n}$ 3d Chern–Simons theory is a topological theory. The Hilbert space ${\mathcal{H}}^{g}$ of this theory placed on a genus-$g$ Riemann surface $\varSigma$ is finite dimensional, with the basis vectors being certain RCFT conformal blocks Witten:1988hf . In the abelian case, there is a continuous variety of boundary conditions, called embeddings in Barbar:2023ncl , and the corresponding blocks are not holomorphic Belov:2005ze ; Aharony:2023zit . A path integral of this theory on $\varSigma$ times an interval $[0,1]$ would calculate a transition amplitude between two states in ${\mathcal{H}}^{g}$. When the theory admits gapped (topological) boundary conditions Kapustin:2010hk ; Kaidi:2021gbs , imposing such a b.c. at one of the boundaries would define a state at the other boundary, the corresponding wavefunction being the partition function of a certain 2d CFT on $\varSigma$.
@@
The bulk Chern–Simons theory coupled to topological gravity by summing over all handlebody topologies gives rise to the Poincaré series, i.e. the sum over modular images of the vacuum character. In case of the genus-one boundary, the vacuum character is the path integral of the CS theory placed on thermal AdS, a handlebody geometry with a particular cycle topologically trivial in the bulk. All other handlebodies are related to thermal AdS by modular transformations. The resulting sum is the Hartle–Hawking wavefunction of “CS gravity.” In fact this definition would call for a sum over all bulk topologies, which is the prescription in the general case Dymarsky:2024frx , but in certain cases such as those discussed in this paper, this sum simplifies to include only handlebodies, or in fact a finite sum over equivalence classes of handlebodies – see section 4.1 for a discussion.
@@ 1.4 Howe duality
The central statement of Howe duality concerns a symplectic group $\mathcal{S}=Sp(2N,F)$ over a field $F$, and “dual pairs,” which are a pair of subgroups of $\mathcal{S}$ which are mutually centralizing and act reductively on a space $V$. The study of dual pairs was introduced by Howe Howe1973Preprint ; Howe1979 ; Howe1989 ; Howe1989a . Howe duality says that a certain representation of $\mathcal{S}$ has a particular structure when restricted to the pair of commuting subgroups. In the present case we will have $F={\mathbb{F}}_{p}$ – the field with $p$ elements ($p$ prime) – and the dual pair is formed by a group $O(n,n,F)$ which acts transitively on the set of codes, and the symplectic group $Sp(2g,F)$ of modular transformations/MacWilliams identities.
@@
For every $g$, each code $\cal C$ has an associated vector $\sum_{c_{i}\in\cal C}\psi_{c_{1},\ldots,c_{g}}$ in this space, which fully determines the genus-$g$ partition function of the corresponding code CFT. Individual terms in the Poincaré series are also vectors in this space. To connect with the picture above, the space $\mathcal{X}$ is isomorphic to the Hilbert space ${\mathcal{H}}^{g}$ of $(U(1)\times U(1))^{n}_{p}$ CS theory on a genus-$g$ Riemann surface, and to the Hilbert space of $2gn$ qudits. The groups $Sp(2g,{\mathbb{F}}_{p})$ and $O^{+}(2n,{\mathbb{F}}_{p})=O(n,n,{\mathbb{F}}_{p})$ both act linearly on $\mathcal{X}$. The group actions commute and $O^{+}(2n,{\mathbb{F}}_{p})\times Sp(2g,{\mathbb{F}}_{p})$ forms a reductive dual pair inside a larger symplectic group $Sp(4ng,{\mathbb{F}}_{p})$, which also acts linearly on $\mathcal{X}$. It is straightforward to show that both the ensemble average and the Poincaré series are invariant under both subgroups.
@@ 2.1.1 $\boldsymbol{(U(1)\times U(1))_{k}}$ Chern–Simons theory on a torus
For simplicity let us start with the level-kkABAB theory:” a single pair of gauge fields $A$ and $B$ on a 3d manifold $\mathcal{M}$ whose boundary is a torus, $\varSigma=\partial\mathcal{M}$. Within our general story this corresponds to the case of $n=g=1$. The 3d action is
@@
The quantization of this theory is reviewed in Aharony:2023zit following Bos:1989wa ; Elitzur:1989nr , see also Belov:2005ze . The first important element is the wavefunction, defined by a path integral on the handlebody (solid torus) with possible Wilson loop insertions, and fixed boundary conditions for the fields $A$ and $B$ at $\partial M$. For level-kkABAB theory” placed on a torus, the Hilbert space $\mathcal{H}$} is a linear combination of $k^{2}$ wavefunctions, known as non-holomorphic blocks in Gukov:2004id ; Belov:2005ze (and codeword blocks in Aharony:2023zit ), given by
@@
The pure phase factor is necessary to make $W[a,b]$ invariant under the shifts of $a,b$ mod $k$, which is confirmed by its action on the wavefunctions (14) Aharony:2023zit ,
@@
Figure 1: The two Wilson loops that wrap a genus-1 surface. The $\gamma_{a}$ loop in (a) is contractible in the bulk so it acts as the identity on the vacuum, while $\gamma_{b}$ in (b) is non-contractible so it acts non-trivially on the vacuum.
@@ 2.1.2 General case: $(U(1)\times U(1))^{n}_{k}$ theory on a genus-$g$ Riemann surface
We now consider the case of a CS theory with $2n$ gauge fields placed on $\mathcal{M}$ with a genus-$g$ boundary $\varSigma=\partial\mathcal{M}$. We choose the standard basis of cycles $\gamma_{\kappa}$, $\kappa=1,2,\dots 2g$ on $\varSigma$ with the symplectic intersection
@@
[…]
@@
We will need the higher-genus generalization of the wavefunctions, forming a basis in the Hilbert space ${\mathcal{H}}^{g}$ of the CS theory on $\varSigma$ of genus $g$. These are defined by $g$-tuples of vectors $(c_{1},\ldots,c_{g})$, $c_{i}\in{\mathbb{Z}}_{k}^{2n}$. From the bulk point of view these are the path integrals on a 3d handlebody geometry $\mathcal{M}$ ending on $\partial M=\varSigma$, with the Wilson line insertions of charges $c_{i}$ wrapping $g$ non-shrinkable cycles. The genus-$g$ handlebody is characterized by a modular parameter $\Omega$, see e.g. Maloney:2020nni . Written explicitly, while taking $\xi=\bar{\xi}=0$, the wavefunctions are
@@
A genus-$g$ wavefunction is a sum over $g$-tuples of vectors $\vec{v}_{I}$, with the index $I$ running from $1,\ldots,g$. In addition, each $\vec{v}_{I}$ is an $2n$-dimensional vector in the lattice $\Lambda^{*}$. In (31), the (implicit) vector indices of $\vec{v}_{I}$ and $\vec{v}_{J}$ are contracted with $\eta$ or with each other. A fixed orthogonal matrix $\mathcal{O}\in O(n,n,\mathbb{R})$ specifies the “embedding” of codes into the Narain moduli space. The sum in (31) is performed over all integer-valued vectors $m_{I}$. The denominator $\Phi(\Omega)$ stands for the one-loop contribution of fluctuations of gauge fields in the bulk, generalizing $|\eta(\tau)|^{2}$ for $g>1$. This factor would be the same for any Chern–Simons theory with the same signature of matrix $K_{ij}$ Belavin:1986cy ; DHoker:1986eaw ; Zograf1987 ; Zograf1987b ; Sarnak1987 ; Krasnov:2000zq ; McIntyre:2004xs ; Gaberdiel:2010jf .
@@
$W[a,b]=\exp\bigg(i\sum_{I,j,j^{\prime}}a_{I}^{j}\,\eta_{jj^{\prime}}\oint_{\gamma^{I}_{a}}A^{j^{\prime}}dx+i\sum_{I,j,j^{\prime}}b_{I}^{j}\,\eta_{jj^{\prime}}\oint_{\gamma^{I}_{b}}A^{j^{\prime}}dx+i{\pi\over k}\sum_{I,j,j^{\prime}}a_{I}^{j}\eta_{jj^{\prime}}b_{I}^{j^{\prime}}\bigg).$ (33)
Here $I$ runs from $1$ to $g$ indexing cycles, and indices $j,j^{\prime}$ run from $1$ to $2n$, indexing gauge fields $A^{j}=\{A_{1},\ldots,A_{n},B_{1},\ldots,B_{n}\}$. As in the case of $g=n=1$, the phase factor in (33) is introduced to make it well defined for $a,b\in{\mathbb{Z}}_{k}^{2n}$.
@@
Generalizing the $g=1$ case of section 2.1.1, the “vacuum” state $|0\rangle:=\Psi_{0,\ldots,0}$ is the CS path integral on the handlebody with shrinkable cycles $\gamma_{a}^{I}$. It can be defined by requiring that any Wilson loop wrapping shrinkable cycles acts trivially,